# CLIX: Cross-Lingual Explanations of Idiomatic Expressions

### v2

@@ 5 Experiments
Models & Metrics. In our experiments, we benchmark the performance of two sequence-to-sequence models, T5 (Raffel et al., 2020) and mT5 (Xue et al., 2021), as well as two LLMs, Llama3.1333Llama3.1 8B Instruct (Dubey et al., 2024) and GPT-3.5 Turbo [^4] Model Version: gpt-3.5-turbo-0125 (OpenAI, 2023). To evaluate our approaches we primarily follow Zhang et al. (2023). To determine the quality of the generated explanations, we calculate their similarity to the gold explanations in the English using a multilingual version of SBERT (Reimers and Gurevych, 2020). We use this metric as it is more likely to assess the meaning of the expressions. Finally, in Tab. 6 we include BLEU and ROUGE scores calculated with our gold translated explanations for the best models using SacreBLEU (Post, 2018) and ROUGE-L.
5.1 Results & Discussion
@@ 6 Human Evaluation
Our annotators found that the model performed extremely well in terms of both Fluency and Accuracy, with average scores of 4.70 and 4.78 respectively. To measure inter-annotator agreement, we calculate Krippendorff’s $\alpha$. These scores range from -1 to 1, where 1 is perfect agreement, 0 is random agreement and -1 is inverse agreement. We found $\alpha=0.642$ for fluency and $\alpha=0.417$ for accuracy. We found two main reasons for disagreement. First, in some cases, an explanation may begin with a literal interpretation and end with a figurative interpretation. In this case, one of our annotators appeared to penalize for the inclusion of the literal interpretation where the other did not. Second, our raters’ native dialects of Spanish are different, resulting in differences in understood meaning or connotation of certain words and phrases.
@@ 7 Conclusion and Future Work
This work represents a first step towards generating cross-lingual explanations of idiomatic expressions. We introduced a dataset consisting of 628 English idioms with their corresponding explanations in English, Spanish and German, and showed that state of the art systems perform well when generating cross-lingual explanations, but not well enough for our ideal target domain of education. For our learned models, we wish to devise more effective strategies that incorporate some notion of directly unwinding figuration and explicit focus on a complete explanation that delivers information so that a learner attains some requisite understanding.
@@
Limitations
Given that we cannot make Oxford-ME public, the primary limitation of our paper is the small size of our publishable dataset. Additionally, we are limited in the number of languages we support. We have only one source language of idiomatic expressions, English. There is a reasonable chance that the models we have experimented with might perform worse in experiments conducted with languages that have smaller amounts of available data. Furthermore, we only attempt cross-lingual explanation into two languages, Spanish and German, which are similarly well supported with data. Finally, the metrics we use in this paper are simply not able to fully capture the phenomenon we wish to measure. We hope to motivate and/or participate in the development of more informative metrics that best suit a task such as ours.
@@ Appendix C Themes
Table 12: EPIE-ME Idiomatic Expression Count ($\geq 5$) per Theme
Theme Num. Idioms
@@
Table 13: Cont’: EPIE-ME Idiomatic Expression Count ($<5$) per Theme
@@ Appendix D Example Prompts
Figure 1: Histograms of Normalized Edit Distance vs Percentage of Data with Bins of Size 0.05 Over EPIE-ME (Spanish & German) and Oxford-ME (Spanish).
@@
Spanish German Spanish
Mean 0.159 0.162 0.069
Min 0 0.092 0
Max 0.781 0.857 0.8
Stdev 0.162 0.216 0.126
Q1 0 0.092 0
Q2 0.131 0.269 0
Q3 0.249 0.440 0.095
Q4 0.781 0.857 0.8
@@ Appendix E Noise in Translations
Where EPIE-ME needed a small amount of correction, Oxford-ME required significantly less modification. [^5] For Oxford-ME, we verify only the Spanish translations for this analysis. We believe this is primarily due to the difference in English explanation length between the two sources of data. Examining the quartiles of normalized edit distance, we see that 75% of the data required roughly less than a 25% modification to the longer string for Spanish explanations in EPIE-ME, less than a 45% modification for German in EPIE-ME, and less than a 10% modification for Spanish explanations in Oxford-ME. Overall, quite low. The high standard deviations can be explained by the lack of samples near the mean, rather a large portion of samples are in the lower quartiles. These measurements can be found in Tab. 15. Graphs of the same information can be found in Fig. 1.
@@ Appendix F Translation as a Method
To test our assumption that translation is not a wholesale solution, we ask GPT-3.5 Turbo to provide only a translation of the idiomatic expressions in our test set. When comparing the semantic similarity of the generated translation to the gold explanation, we see that this approach appears to perform better than that of the fine-tuned models used in our experiments in Tab 3, earning scores of 45.80 and 44.21 for Spanish and German respectively. However, we asked our annotators to take a closer look at the outputs to make note of 3 things for each sample; Is the translated expression good/natural? Is the translation done word-for-word? Is this translation an instance of an idiom in your native language?

### v3

@@ 5 Experiments
Models and Metrics
In our experiments, we benchmark the performance of two sequence-to-sequence models, T5 (Raffel et al., 2020) and mT5 (Xue et al., 2021), as well as two LLMs, Llama3.1333Llama3.1 8B Instruct (Dubey et al., 2024) and GPT-3.5 Turbo [^4] Model Version: gpt-3.5-turbo-0125 (OpenAI, 2023). To evaluate our approaches we primarily follow Zhang et al. (2023). To determine the quality of the generated explanations, we calculate their similarity to the gold explanations in the English using a multilingual version of SBERT (Reimers and Gurevych, 2020). We use this metric as it is more likely to assess the meaning of the expressions. Finally, in Tab. 6 we include BLEU and ROUGE scores calculated with our gold translated explanations for the best models using SacreBLEU (Post, 2018) and ROUGE-L.
5.1 Results and Discussion
@@ 6 Human Evaluation
Our annotators found that the model performed extremely well in terms of both Fluency and Accuracy, with average scores of 4.70 and 4.78 respectively. To measure inter-annotator agreement, we calculate Krippendorff’s $\alpha$. These scores range from -1 to 1, where 1 is perfect agreement, 0 is random agreement and -1 is inverse agreement. We found $\alpha=0.642$ for fluency and $\alpha=0.417$ for accuracy. We found two main reasons for disagreement. First, in some cases, an explanation may begin with a literal interpretation and end with a figurative interpretation. In this case, one of our annotators appeared to penalize for the inclusion of the literal interpretation where the other did not. Second, our raters’ native variants of Spanish are different, resulting in differences in understood meaning or connotation of certain words and phrases. However, we do not consider difference in variant a significant issue as it pertains to our overall results. Since the generative models examined in this paper are trained on the same generalized Spanish data that the embedding model is trained on (all ignoring the existence of any variants), the scores given by our annotators ought to be lower than an automated evaluation rather than higher (assuming the automated evaluation is infallible).
@@ 7 Conclusion and Future Work
This work represents a first step towards generating cross-lingual explanations of idiomatic expressions. We introduced two datasets, one consisting of 628 English idioms and another containing 6218 in English idioms, each with their explanations in English, Spanish and German, and showed that state of the art systems perform well when generating cross-lingual explanations, but not well enough in certain aspects for our ideal target domain of education. For our learned models, we wish to devise more effective strategies that incorporate some notion of directly unwinding figuration and explicit focus on a complete explanation that delivers information so that a learner attains some requisite understanding.
@@
8 Limitations
Given that we cannot make Oxford-ME public, the primary limitation of our paper is the small size of our publishable dataset. However, we are able to provide this data to other researchers upon request. Additionally, we are limited in the number of languages we support. We have only one source language of idiomatic expressions, English. There is a reasonable chance that the models we have experimented with might perform worse in experiments conducted with languages that have smaller amounts of available data. Furthermore, we only attempt cross-lingual explanation into two languages, Spanish and German, which are similarly well supported with data. Finally, the metrics we use in this paper are simply not able to fully capture the phenomenon we wish to measure. We hope to motivate and/or participate in the development of more informative metrics that best suit a task such as ours.
Acknowledgments
We thank members of the BLAST and NALA labs for their feedback. This work utilized the Blanca condo computing resource at the University of Colorado Boulder. Blanca is jointly funded by computing users and the University of Colorado Boulder.
@@ Appendix C Themes
Table 12: EPIE-ME Idiomatic Expression Count ($\geq 10$) per Theme
Theme Num. Idioms
@@
Table 13: Cont’: EPIE-ME Idiomatic Expression Count ($<10$) per Theme
@@ Appendix D Example Prompts
Figure 1: Histograms of Normalized Edit Distance vs Percentage of Data with Bins of Size 0.05 Over Both Languages in EPIE-ME and Oxford-ME.
@@
Spanish German Spanish German
Mean 0.159 0.283 0.131 0.183
Min 0 0.092 0 0
Max 0.781 0.857 0.8 0.889
Stdev 0.162 0.216 0.165 0.241
Q1 0 0.092 0 0
Q2 0.131 0.269 0.071 0.043
Q3 0.249 0.440 0.205 0.307
Q4 0.781 0.857 0.839 .89
@@ Appendix E Noise in Translations
Where EPIE-ME needed a small amount of correction, Oxford-ME required significantly less modification. [^5] For Oxford-ME, we verify only the Spanish translations for this analysis. We believe this is primarily due to the difference in English explanation length between the two sources of data. Examining the quartiles of normalized edit distance, we see that 50% of the data required roughly less than a 14% modification to the Spanish explanations in EPIE-ME, less than a 27% modification for German in EPIE-ME, less than an 8% modification for Spanish explanations in Oxford-ME, and less than a 4% modification for German in Oxford-ME. Overall, quite low. The high standard deviations can be explained by the lack of samples near the mean, rather a large portion of samples are in the lower quartiles. These measurements can be found in Tab. 15. Graphs of the same information can be found in Fig. 1.
@@ Appendix F Translation as a Method
To test our assumption that translation is not a wholesale solution, we ask GPT-3.5 Turbo to provide only a translation of the idiomatic expressions in our test set. When comparing the semantic similarity of the generated translation to the gold explanation, we see that this approach appears to perform better than that of the fine-tuned models used in our experiments in Tab 3, earning scores of 45.80 and 44.21 for Spanish and German respectively. However, we asked our annotators to take a closer look at the outputs to make note of 3 things for each sample: Is the translated expression good/natural? Is the translation done word-for-word? Is this translation an instance of an idiom in your native language?